c)You have 0.6 mL an undiluted culture at a density of 3.7 x 107 CFU/mL. You add 5.4 mL of sterile diluent. What is the dilution and final cell density?

e) Plating 2.0 mL of a sample diluted by a factor or 10-2 procedures 56 colonies. What was the original cell density in the sample?

f)Plating 0.2 mL of a sample diluted by a factor or 10-4 procedures 87 colonies. What was the original cell density in the sample?
Answer #1

c) dilution = aliquout volume/final volume

aliquot volume= 0.6 mL final volume= 5.4+0.6=6 mL

dilution =0.6/6= 0.1

dilution factor= 6/0.6=10

final cell density= concentration of cells in the undiluted sample* dilution

= 3.7 x 107 CFU/mL *0.1

= 3.7*10^6 CFU/mL

e) CFU/mL= number of colonies/( volume plated in mL * dilution)

number of colonies= 56, volume plated= 2 mL dilution=10^-2

CFU/mL= 56/(2*10^-2)

= 2800

CFU/mL= 2.8*10^3 CFU/mL

f)

CFU/mL= number of colonies/( volume plated in mL * dilution)

number of colonies= 87, volume plated=0.2 mL dilution=10^-4

CFU/mL= 87/ (10^-4*0.2)

=4.35*10^6 CFU/mL

so cell density = 4.35*10^6 CFU/mL